# Beyond Accuracy: A Stability-Aware Metric for Multi-Horizon Forecasting

### v1

@@ Abstract.
Traditional time series forecasting methods optimize for accuracy alone. This objective neglects temporal consistency, in other words, how consistently a model predicts the same future event as the forecast origin changes. We introduce the forecast accuracy and coherence score (forecast AC score for short) for measuring the quality of probabilistic multi-horizon forecasts in a way that accounts for both multi-horizon accuracy and stability. Our score additionally provides for user-specified weights to balance accuracy and consistency requirements. As an example application, we implement the score as a differentiable objective function for training seasonal ARIMA models and evaluate it on the M4 Hourly benchmark dataset. Results demonstrate substantial improvements over traditional maximum likelihood estimation. Our AC-optimized models achieve a 75% reduction in forecast volatility for the same target timestamps while maintaining comparable or improved point forecast accuracy.
@@ 3.1. Problem definition
Consider a multi-horizon forecasting scenario where the forecaster receives new information at each timestamp $t$. The information available to the forecaster at time $t$ is denoted as $\mathcal{Z}_{t}=(Z_{1},Z_{2},\cdots,Z_{t})$. $Z_{t}$ represents the new information that arrives at $t$. The forecaster then generates $m$-step-ahead forecasts using all relevant information available up to (and including) time $t$. The $m$-step-ahead forecasts issued at origin time $t$ are denoted as $\hat{y}_{t+1|t},\hat{y}_{t+2|t},\cdots,\hat{y}_{t+m|t}$, where the subscripts $t+j$ correspond to the forecast target time and $t$ corresponds to the forecast origin. We assume the $m$-step-ahead forecasts are sampled from a joint probability distribution $f_{t}(\hat{y}_{t+1|t},\hat{y}_{t+2|t},\cdots,\hat{y}_{t+m|t})$.
Remark.
In situations where there are no covariates, the new information $Z_{t}$ is simply equal to the newly observed target value $y_{t}$. In the presence of covariates, $Z_{t}=(x_{t},y_{t})$ where $x_{t}$ are all covariates which are used to generate the forecasts.
@@
are forecasts targeting the time $t$ issued at different origins $t-m,t-m+1,\cdots,t-1$, respectively. Each anti-diagonal therefore collects the sequence of forecast revisions for a particular target time.
@@ 3.2. Measuring the distance between forecast distributions
where $\|(x_{1},\cdots,x_{n})\|=(\sum x_{i}^{2})^{\frac{1}{2}}$ is the $L^{2}$ norm. Then we recover the squared Energy Distance [8]. Substituting the Dirac distribution supported at the true outcome $y$ into $G$, we recover the Energy Score [3]
@@
(13) $\mathrm{Stb}(F|\mathcal{Z})=\frac{1}{n}\sum_{t=1}^{n}D^{2}(\operatorname*{marg}_{\hat{Y}_{t+1|t}}f_{t},\operatorname*{marg}_{\hat{Y}_{t+m+1|t+1}}f_{t+1}).$
@@ 3.5. The Choice of Weights
For a fixed target time $T$ and the earliest origin denoted as $i_{\min}$, let the decision time $\tau$ be a time between the earliest origin and the target time, i.e., $i_{\min}\leq\tau<T$. Then, it is convenient to define the distance to the the decision time from the origin as $u=|i-\tau|$ and to define a corresponding weighting function $\omega(u)$. Similarly, define the horizon $h$ as the distance from origin to target time, h=Tih=T-i, and the corresponding weighting function $\omega(h)$. Depending upon the operational context, one representation may be more natural than the other, i.e., a representation based on distance from the origin to the decision time versus a representation based on distance from the origin to the target time. Below, we will develop the case for $\omega(h)$; that for $\omega(u)$ is analogous.
@@
    Normalization: $\sum_{h}\omega(h)=1$. Ensures compatibility across horizons and across datasets.
@@
    Monotonicity toward the target time (optional): weights increase as we approach the decision point, $\omega(h_{1})\leq\omega(h_{2})$ whenever $h_{1}>h_{2}$. It is assumed that later forecasts are more consequential. However, the framework allows for alternative shapes (e.g., U-shaped) depending upon domain needs.
@@
    Hyperbolic: $\omega(h)=\frac{1}{1+\beta h}$
@@
    Inverse-variance: $\omega(h)=\frac{1}{\mathrm{Var}(f_{h})}$
@@
Accuracy and stability capture different aspects of forecast quality. Accuracy measures closeness to the true target, while stability measures temporal consistency across origins. Their operational relevance may differ: late revisions can have high costs, affecting stability, while accuracy may be more relevant across all horizons. The proposed framework allows the use of separate weighting functions, $\omega(h)$ for accuracy and $\omega^{\prime}(h)$ for stability, to reflect these differences. In applications where the same importance applies, the same function can be used for both.
@@ 4. Experiments
Most implemented machine learning forecasting models employ one of these standard metrics as their training objective. However, a significant limitation of current practice is that loss evaluation typically focuses on one-step-ahead forecast accuracy. This limitation is suboptimal in certain situations. On one hand, optimizing for one step or a fixed horizon often results in models that perform well at the specific horizons but deteriorate at other horizons. In fact, many practical applications require evaluating forecast quality across the entire prediction window rather than at isolated time points. For example, energy system operators may require reliable load forecasts from 1 hour ahead up to 7 days ahead simultaneously in order to make operational decisions. A loss function that treats each horizon independently can potentially lead to suboptimal real-world performance despite strong performance on standard benchmarks.
@@
In this section, we show how to set up a seasonal ARIMA model with our metric as the objective function. The trained model is compared with its counterpart trained using traditional objectives. Out-of-sample tests are performed, in which we see that models trained using our metric achieve better accuracy and stability over long horizons.
@@ 4.1. Dataset
We evaluate our method on the M4 Hourly dataset [6], a widely-used forecasting benchmark consisting of 414 hourly time series from diverse domains. This dataset was introduced in the M4 forecasting competition and has become a standard benchmark dataset for evaluating forecasting methods. For computational efficiency, we randomly selected a subset of 50 time series from the full M4 Hourly dataset for our experiments. The selection was performed using a fixed random seed (seed=42) to ensure reproducibility of results.
@@ 4.2. Model explanation
We employ the seasonal autoregressive integrated moving average (SARIMA) model. SARIMA captures both non-seasonal and seasonal patterns in the data. Most of the time series in the M4 dataset exhibits seasonal behavior. Thus SARIMA models serve as suitable baseline models for demonstrating the performance of our metric.
@@ A SARIMA model is specified by seven hyperparameters: SARIMA($p$,$d$,qq)×($P$,$D$,$Q$,$s$) where:
The mathematical representation of the SARIMA model is given by
(18) $\Phi(L^{s})\phi(L)(1-L^{s})^{D}(1-L)^{d}y_{t}=\Theta(L^{s})\theta(L)\epsilon_{t}$
@@
$\theta(x)=1-\theta_{1}x-\cdots\theta_{q}x^{q},$
$\Theta(x)=1-\Theta_{1}x-\cdots\Theta_{Q}x^{Q}.$
For a fixed set of hyperparameters $(p,d,q,P,D,Q,s)$, the model is determined by the corresponding autoregressive and seasonal autoregressive coefficients $\phi_{j},\Phi_{j},\theta_{j},\Theta_{j}$. This simple structure allows us to easily set up optimization procedures to find the optimal coefficients under our metric.
@@ 4.3. Training
Hyperparameter Selection.We employ the auto-ARIMA algorithm from the pmdarima library [10] to automatically determine the optimal SARIMA hyperparameters $(p,d,q,P,D,Q,s)$. The auto-ARIMA performs a stepwise search over the hyperparameter space. The Akaike Information Criterion (AIC) is used as the metric for comparison, which accounts for both model fit and model complexity. This automated hyperparameter selection procedure ensures that our models are tuned to the specific pattern of each time series in the dataset.
Traditional Model Fitting. Using the hyperparameters found by auto-ARIMA, we create a traditional SARIMA model using statsmodel library [9]. The model is fitted on the training set using maximum likelihood estimation (MLE) with standard one-step-ahead forecast error minimization as the optimization objective. This serves as a benchmark for comparing with the optimized model under our novel metric.
Differentiable Model Fitting. We start by creating a differentiable SARIMA model with the same hyperparameters as the traditional model. The autoregressive coefficients are treated as learnable parameters. They are initialized as random floats in the interval (1,1)(-1,1).
Our innovation lies in using our multi-horizon stability-aware forecast AC score as the objective for learning the autoregressive coefficients. In each training epoch, we generate rolling forecasts on the training data. Specifically, the rolling forecast starts from the timestamp whose index is equal to the minimum required history length, which is in turn equal to the maximum lag of the SARIMA model. We produce $H$-step-ahead forecasts for each valid forecast origin. The multi-step forecasts are simulated via the equation (18). The forecast ensembles are stored in the form of (2). The forecast origins are processed in batches of size 32 for computational efficiency.
@@
Since the setup of the differentiable model as well as the entire forecast generating process are implemented with PyTorch, the whole process is differentiable. We backpropagate to compute the gradient of the loss function $\nabla_{\phi,\Phi,\theta,\Theta}S_{A}C(F|Z)$ where $\phi,\Phi,\theta,\Theta$ refer to all of the autoregressive coefficients. We employed the AdamW optimizer with an initial learning rate of 0.05. A ReduceLROnPlateau scheduler was employed to reduce the learning rate by a factor of 0.5 when the loss plateaus for 10 consecutive epochs.
After the training is completed, we verify that the SARIMA model with the learned coefficients are stationary by examining the roots of the autoregressive polynomial
@@
Out-of-sample Testing. In order to validate the results on unseen data, we perform out-of-sample tests. We use both of the traditionally trained model and the model trained with our metric to generate rolling forecasts on the test set, for each time series in the dataset.
@@ 4.4. Results
We evaluated the performance of our AC-optimized SARIMA model against the traditionally trained model on 50 time series in the M4 dataset. Overall, results show that our custom optimization yields substantially improved stability, while often with minimal accuracy trade-offs.
Overall Performance We measure the out-of-sample forecasts of our AC-optimized model using our forecast AC score (14). Out of 50 test series, 45 demonstrated better scores. The median of the improvement in the stability score is 55%, with slightly improved multi-horizon accuracy score by 4.8%.
@@
Enhanced Forecast Stability. The ultimate purpose of our metric is to guide the model to output more stable forecasts. This has been achieved in the out-of-sample tests. Figure 2 displays forecasts generated from multiple forecast origins, all targeting the same future timestamp. Traditionally trained SARIMA models exhibit substantial variation across these forecasts. For example, in Figure 2(a) the traditionally trained model updates the forecast by considerable amounts when moving through forecast origins, and often completely reverts its previous update after another step. In contrast, our AC-optimized model produces remarkably consistent predictions with forecasts issued from different origins converging to similar values as the traditionally trained model predicts when approaching target timestamps.
(a) Vertical Forecasts Targeting t = 90
(b) Vertical Forecasts Targeting t = 150
(c) Vertical Forecasts Targeting t = 220
(d) Vertical Forecasts Targeting t = 290
Figure 2. Vertical Forecasts of Unit T000358 in M4 Dataset
Accuracy Stability Trade-off. A natural concern when optimizing for stability is whether this comes at the cost of forecast accuracy. Our empirical results indicate that this trade-off is remarkably favorable. For 90.0% of the time series, we achieved substantial improvements in forecast stability, with the stability score (13) improving by a median of 55% while maintaining comparable or even improved point forecast accuracy. In fact, the discounted multi-horizon accuracy score (15) was slightly improved for most of the test series. The median of this percentage improvement in the accuracy score was 4.8%.
@@
To quantify the improvement in forecast stability in another way, we compute the standard deviation of forecasts targeting the same future timestamp across different forecast origins. Specifically, for each target timestamp $t$ in the test set, we collect all forecasts $\hat{y}_{t|s}$, $t-H\leq s\leq t-1$ and calculate their standard deviation
$\sigma_{\text{vertical}}(t)=\mathrm{std}_{s}(\hat{y}_{t|s}).$
@@
We then average over all target timestamps $t$. This mean value measures the volatility of vertical forecasts. Lower values indicate that predictions for the same future point are stable throughout forecast origins. We then average this mean standard deviation across all test time series to obtain an overall stability measurement. As a result, our AC-optimized model achieved a average standard deviation of $2.89$, compared to $11.76$ for the traditionally trained model. This represents a relative reduction of 75.4% in forecast volatility for the same target timestamps.

### v2

@@ Abstract.
Traditional time series forecasting methods optimize for accuracy alone. This objective neglects temporal consistency, in other words, how consistently a model predicts the same future event as the forecast origin changes. We introduce the forecast accuracy and coherence score (forecast AC score for short) for measuring the quality of probabilistic multi-horizon forecasts in a way that accounts for both multi-horizon accuracy and stability. Our score additionally allows user-specified weights to balance accuracy and consistency requirements. As an example application, we implement the score as a differentiable objective function for training seasonal auto-regressive integrated models and evaluate it on the M4 Hourly benchmark dataset. Results demonstrate substantial improvements over traditional maximum likelihood estimation. Regarding stability, the AC-optimized model generated out-of-sample forecasts with 91.1% reduced vertical variance relative to the MLE-fitted model. In terms of accuracy, the AC-optimized model achieved considerable improvements for medium-to-long-horizon forecasts. While one-step-ahead forecasts exhibited a 7.5% increase in MAPE, all subsequent horizons experienced an improved accuracy as measured by MAPE of up to 26%. These results indicate that our metric successfully trains models to produce more stable and accurate multi-step forecasts in exchange for some degradation in one-step-ahead performance.
@@ 3.1. Problem definition
Consider a multi-horizon forecasting scenario where the forecaster receives new information at each timestamp $t$. The information available to the forecaster at time $t$ is denoted as $\mathcal{Z}_{t}=(Z_{1},Z_{2},\cdots,Z_{t})$. $Z_{t}$ represents the new information that arrives at $t$, which contains the newly realized value of the target time series and any covariates that become known at time t. The forecaster then generates $m$-step-ahead forecasts using all relevant information available up to (and including) time $t$. The $m$-step-ahead forecasts issued at origin time $t$ are denoted as $\hat{y}_{t+1|t},\hat{y}_{t+2|t},\cdots,\hat{y}_{t+m|t}$, where the subscripts $t+j$ correspond to the forecast target time and $t$ corresponds to the forecast origin. We assume the $m$-step-ahead forecasts are sampled from a joint probability distribution $f_{t}(\hat{y}_{t+1|t},\hat{y}_{t+2|t},\cdots,\hat{y}_{t+m|t})$.
@@
are forecasts targeting the time $t$ issued at different origins $t-m,t-m+1,\cdots,t-1$, respectively. Each anti-diagonal therefore collects the sequence of forecast revisions for a particular target time. Vertical stability refers to the volatility of the sequence (3), measured by its variance averaged over t.
@@ 3.2. Measuring the distance between forecast distributions
where $\|(x_{1},\cdots,x_{n})\|=(\sum x_{i}^{2})^{\frac{1}{2}}$ is the $L^{2}$ norm. Then we recover the squared Energy Distance [7]. Suppose the random variable $Y\sim G$ is realized and the realized value is $y$. One may view the realization as a Dirac measure supported at y. Substituting it into $G$, we recover the Energy Score [2]
@@
(13) $\mathrm{Stb}(F|\mathcal{Z})=\frac{1}{n-1}\sum_{t=1}^{n-1}D^{2}(\operatorname*{marg}_{\hat{Y}_{t+1|t}}f_{t},\operatorname*{marg}_{\hat{Y}_{t+m+1|t+1}}f_{t+1}).$
@@
Remark.
The metric evaluates forecast quality based solely on the resulting distributions, without assuming how those forecasts are produced. This can be applied to diverse types of multi-horizon forecasting models, for example
  • Iterative forecasting: Multi-step forecasts generated by repeatedly applying one-step-ahead models, common in classical time series methods and transformer-based architectures.
  • Direct multi-horizon forecasting: Models trained to produce all forecast horizons simultaneously, which is commonly used in modern neural architectures.
  • Skeleton-based forecasting: Generating forecasts at sparse key horizons (e.g., $h\in\{24,48,72\}$ for hourly data) and using interpolation, extrapolation, or learned smoothing to obtain intermediate forecasts.
  • Attention-based architectures: Transformer models like Temporal Fusion Transformer that use self-attention mechanisms to capture temporal dependencies and generate multi-horizon forecasts.
@@ 3.3. Empirical Approximation
Remark (Point Forecast as a Special Case).
The forecast AC score can be extended to evaluating point forecasts. In fact, the point estimate $\hat{y}$ may be viewed as a Dirac distribution supported at $\hat{y}$. Thus the empirical formula of accuracy score (15) becomes
(17) $Acc(F|Z)=\frac{1}{n}\sum_{t=1}^{n}\left(\sum_{j=1}^{m}w_{j}|\hat{y}_{t+j|t}-y_{t+j}|^{2}\right)^{\frac{1}{2}}$
and (16) becomes
(18) $Stb(F|Z)=\frac{1}{n-1}\sum_{t=1}^{n-1}\left(\sum_{j=2}^{m}w_{j}|\hat{y}_{t+j|t}-\hat{y}_{t+j|t+1}|^{2}\right)^{\frac{1}{2}}.$
@@ 3.5. The Choice of Weights
For a fixed target time $T$ and the earliest origin denoted as $i_{\min}$, let the decision time $\tau$ be a time between the earliest origin and the target time, i.e., $i_{\min}\leq\tau<T$. Then, it is convenient to define the distance to the the decision time from the origin as $u=|i-\tau|$ and to define a corresponding weighting function $w(u)$. Similarly, define the horizon $h$ as the distance from origin to target time, h=Tih=T-i, and the corresponding weighting function $w(h)$. Depending upon the operational context, one representation may be more natural than the other, i.e., a representation based on distance from the origin to the decision time versus a representation based on distance from the origin to the target time. Below, we will develop the case for $w(h)$; that for $w(u)$ is analogous.
@@
    Normalization: $\sum_{h}w(h)=1$. Ensures compatibility across horizons and across datasets.
@@
    Monotonicity toward the target time (optional): weights increase as we approach the decision point, $w(h_{1})\leq w(h_{2})$ whenever $h_{1}>h_{2}$. It is assumed that later forecasts are more consequential. However, the framework allows for alternative shapes (e.g., U-shaped) depending upon domain needs.
@@
    Hyperbolic: $w(h)=\frac{1}{1+\beta h}$
@@
    Inverse-variance: $w(h)=\frac{1}{\mathrm{Var}(f_{h})}$
@@
Accuracy and stability capture different aspects of forecast quality. Accuracy measures closeness to the true target, while stability measures temporal consistency across origins. Their operational relevance may differ: late revisions can have high costs, affecting stability, while accuracy may be more relevant across all horizons. The proposed framework allows the use of separate weighting functions, $w(h)$ for accuracy and $w^{\prime}(h)$ for stability, to reflect these differences. In applications where the same importance applies, the same function can be used for both.
@@ 4. Experiments
Most implemented machine learning forecasting models employ one of these standard metrics as their training objective. However, a significant limitation of current practice is that loss evaluation typically focuses on one-step-ahead forecast accuracy. This limitation is suboptimal in certain situations, as optimizing for one step or a fixed horizon often results in a model that performs well at the specific horizon but deteriorates at other horizons. In fact, many practical applications require evaluating forecast quality across the entire prediction window rather than at isolated time points. For example, energy system operators may require reliable load forecasts from 1 hour ahead up to 7 days ahead simultaneously in order to make operational decisions. A loss function that treats each horizon independently can potentially lead to suboptimal real-world performance despite strong performance on standard benchmarks.
@@
In this section, we show how to set up a seasonal auto-regressive integrated model with our metric as the objective function. The trained model is compared with its counterpart trained using traditional objectives. Out-of-sample tests are performed, in which we see that models trained using our metric achieve better accuracy and stability over long horizons.
@@ 4.1. Dataset
We evaluate our method on the M4 Hourly dataset [5], a widely-used forecasting benchmark consisting of 414 hourly time series from diverse domains. This dataset was introduced in the M4 forecasting competition and has become a standard benchmark dataset for evaluating forecasting methods.
@@ 4.2. Model explanation
We employ the seasonal autoregressive integrated moving average model (SARIMA). SARIMA captures both non-seasonal and seasonal patterns in the data. Most of the time series in the M4 dataset exhibit seasonal behavior. Thus, SARIMA models serve as suitable baseline models for demonstrating the performance of our metric.
@@ A SARIMA model is specified by seven hyperparameters: SARIMA($p$,$d$,qq)×($P$,$D$,$Q$,$s$) where:
For computational simplicity, we restrict the moving average order $q$ and $Q$ to be 0. In other words, our model takes the form of SARIMA(p,d,0)x(P,D,0,s). We will refer to this model as SARI in the paper. The mathematical representation of our model is given by
(20) $\Phi(L^{s})\phi(L)(1-L^{s})^{D}(1-L)^{d}y_{t}=\epsilon_{t}$
@@
For a fixed set of hyperparameters $(p,d,P,D,s)$, the model is determined by the corresponding autoregressive and seasonal autoregressive coefficients $\phi_{j},\Phi_{j}$. This simple structure allows us to easily set up optimization procedures to find the optimal coefficients under our metric.
@@ 4.3. Training
Hyperparameter Selection. We employ the auto-ARIMA algorithm from the pmdarima library [9] to automatically determine the optimal SARI hyperparameters $(p,d,P,D,s)$. The auto-ARIMA performs a stepwise search over the hyperparameter space. The Akaike Information Criterion (AIC) is used as the metric for comparison, which accounts for both model fit and model complexity. This automated hyperparameter selection procedure ensures that our models are tuned to the specific pattern of each time series in the dataset.
Traditional Model Fitting. Using the hyperparameters found by auto-ARIMA, we create a traditional SARI model using statsmodel library [8]. The model is fitted on the training set using maximum likelihood estimation (MLE) with standard one-step-ahead forecast error minimization as the optimization objective. This serves as a benchmark for comparing with the optimized model under our novel metric.
Differentiable Model Fitting. We start by creating a differentiable SARI model with the same hyperparameters as the traditional model. The autoregressive coefficients are treated as learnable parameters. They are initialized as random floats in the interval (1,1)(-1,1).
Our innovation lies in using our multi-horizon stability-aware forecast AC score as the objective for learning the autoregressive coefficients. In each training epoch, we generate rolling forecasts on the training data. Specifically, the rolling forecast starts from the timestamp whose index is equal to the minimum required history length, which is in turn equal to the maximum lag of the SARI model. We produce $H$-step-ahead forecasts for each valid forecast origin. The multi-step forecasts are simulated via the equation (20). The forecast ensembles are stored in the form of (2). The forecast origins are processed in batches of size 32 for computational efficiency.
@@
Since the setup of the differentiable model as well as the entire forecast generating process are implemented with PyTorch, the whole process is differentiable. We backpropagate to compute the gradient of the loss function $\nabla_{\phi,\Phi}S_{AC}(F|Z)$ where $\phi,\Phi$ refer to all of the autoregressive coefficients. We employed the AdamW optimizer with an initial learning rate of 0.05. A ReduceLROnPlateau scheduler was employed to reduce the learning rate by a factor of 0.5 when the loss plateaus for 10 consecutive epochs. We set the seed for PyTorch and Numpy operations to 42 for reproducibility.
After the training is completed, we verify that the SARI model with the learned coefficients are stationary by examining the roots of the autoregressive polynomial
@@
Out-of-sample Testing. In order to validate the results on unseen data, we perform out-of-sample tests. We use both of the traditionally trained model and the model trained with our metric to generate rolling forecasts on the test set, for each time series in the dataset. The models we used in the experiments generate and evaluate point forecasts, which can be viewed as degenerate probability distributions with all mass concentrated at a single value. With this simplification we can demonstrate the metric’s effectiveness in improving multi-horizon accuracy and stability without the additional complexity of generating full predictive distributions.
@@ 4.4. Results
We evaluated the performance of our AC-optimized SARI model against the traditionally trained model on 414 time series in the M4 dataset. Overall, results show that our custom optimization yields substantially improved stability, while often with minimal accuracy trade-offs, whether measured by one-step-ahead error (using MAPE) or multi-step-ahead error and stability (using forecast AC Score). Figure 2 provides a comprehensive comparison of the MLE-fitted and AC-optimized models in terms of their forecast AC score (14), accuracy score (9) and stability score (13). Specifically, figures 2(a), 2(c), and 2(e) display these log transformed version of these scores at each percentile (log transformation is applied due to the large variation in original scales across time series). Figures 2(b), 2(d), and 2(f) show the percentile distribution of relative improvements achieved by the AC-optimized model over the MLE-fitted model for each of the three metrics. We quote some statistics of the relative improvement in Table 1.
Table 1. Relative improvement of AC-optimized model over MLE-fitted model
Statistic AC Score Improvement Accuracy Score Improvement Stability Score Improvement
25% Percentile 2.38% -2.08% 27.06%
50% Percentile (Median) 7.94% 3.18% 50.00%
75% Percentile 17.72% 11.09% 81.00%
(a) Log AC Score
(b) Relative Improvement of AC Score
(c) Log Accuracy Score
(d) Relative Improvement of Accuracy Score
(e) Log Stability Score
(f) Relative Improvement of Stability Score
Figure 2. Comparison of metrics at log scale
@@
Enhanced Forecast Stability. The ultimate purpose of our metric is to guide the model to output more stable forecasts. This has been achieved in the out-of-sample tests. Figure 3 displays forecasts generated from multiple forecast origins, all targeting the same future timestamp. The traditionally trained SARI model exhibits substantial variation across these forecasts. It updates the forecast by considerable amounts when moving through forecast origins, and often completely reverts its previous update after another step. In contrast, our AC-optimized model produces remarkably consistent predictions with forecasts issued from different origins converging to similar values as the traditionally trained model predicts when approaching target timestamps. Figure 3 visualizes such scenarios where the x-axis represents the forecast origins and the y-axis represents the forecast values made at these origins targeting the same target timestamp.
Figure 3. Examples illustrating superior vertical stability.
@@
To quantify the improvement in forecast stability in an alternative way, we compute the sample variance of the forecasts targeting the same future timestamp across different forecast origins. Specifically, for each target timestamp $t$ in the test set, we collect all forecasts $\hat{y}_{t|s}$, $t-h\leq s\leq t-1$ and calculate their variance
(22) $s^{2}_{\text{vertical}}(t)=\frac{1}{h-1}\sum_{s=t-h}^{t-1}(\hat{y}_{t|s}-\bar{\hat{y}})^{2}$
@@
where $\bar{\hat{y}}$ is the average of forecasts $\hat{y}_{t|s}$ over $t-h\leq s\leq t-1$. We then average over all target timestamps $t$. This mean value measures the volatility of vertical forecasts. Lower values indicate that predictions for the same future point are stable throughout forecast origins. We then average this mean variance across all test time series to obtain an overall stability measurement. As a result, the vertical volatility (represented by (22)) of the AC-optimized model is only 8.9% of that of the MLE-fitted model, with the former output forecasts whose variance is $4.1$ and the latter’s forecasts having a much greater variance equal to $45.7$.
Accuracy Trade-off at Different Horizons. Two natural concerns when using the forecast AC score are
  1. (1)
    Whether improving stability comes at the cost of forecast accuracy.
  2. (2)
    Whether aiming at multi-step-ahead comes at the cost of accuracy at one-step-ahead.
To evaluate these trade-offs comprehensively, we employ two complementary evaluation strategies. We begin by assessing performance using our forecast AC score as it directly reflects the optimization objective and both the accuracy and stability of multi-step ahead. We then supplement this with horizon-specific MAPE analysis for two reasons: (1) MAPE provides an intuitive, scale-independent accuracy measure familiar to most practitioners and allows for more straightforward interpretation; (2) examining individual horizons allows us to demonstrate that our multi-horizon optimization does not drastically sacrifice one-step-ahead forecast quality, which is still an important objective for most forecasting works.
The AC-optimized model achieved improved forecast AC score for 83% of the tested series, with significantly improved stability as shown in Figure 2(f). In exchange for the stability improvement, the AC-optimized model generated slightly less accurate forecasts for approximately 25% of the time series (the first quantile in Figure 2(d)). However, the combined AC score still improved for the majority of series, demonstrating that stability gains outweighed the modest accuracy losses for these cases.
Next we measure the trade-off between the one-step-ahead accuracy and multi-step-ahead accuracy plus stability by evaluating the MAPE of the one-step-ahead forecasts. Figure 4 presents the percentage MAPE improvement of the AC-optimized model relative to the traditional MLE-fitted model across forecast horizons, along with 50% confidence intervals. The comparison is evaluated with the percentage improvement in terms of MAPE, as in equation (23)
(23) % $\text{\% Improvement}=\frac{\text{MAPE}_{MLE}-\text{MAPE}_{AC}}{\text{MAPE}_{MLE}}.$
At the one-step-ahead horizon, our AC-optimized model produces moderately less accurate forecasts, with its MAPE 7.5% worse compared to that of the traditional model at 50% percentile. However, this short-term accuracy cost is quickly recovered: starting at a horizon equal to 2, our model consistently outperforms the traditional approach in both accuracy and vertical stability. The performance gap widens substantially as the forecast horizon extends. For longer horizons, our AC-optimized model achieves up to a 26% better MAPE compared to the traditional model. This pattern demonstrates that, while traditional MLE training excels at optimizing one-step-ahead predictions, it fails to generalize effectively to longer horizons. In contrast, our multi-horizon approach is able to generate much more accurate long-horizon forecasts with only a relatively small loss of one-step-ahead forecast accuracy. This trade-off can prove useful for practical decision-making where medium and long term planning decisions are typically important.
Figure 4. Percentage MAPE improvement of the AC-optimized model over the traditional MLE-fitted model.
Remark.
One may notice that (23) starts to decrease after horizon equal to 12. This is due to our choice of horizon discounting weights, which is linear in horizon and assigns more importance to shorter horizon compared to longer horizon. The improvement remains positive, indicating that the AC-optimized model still outperforms the MLE-fitted model to a lesser extent at these horizons.
One-Step-Ahead Accuracy Distribution. At one step ahead, the AC-optimized model trades off accuracy to some extent in exchange for longer horizon accuracy and stability. To provide a more detailed picture of this trade-off, we analyze the distribution of performance differences between the two approaches. Table 2 provides detailed statistics that compare the performance of two models measured by the one-step-ahead MAPE. Figure 5 is a histogram of the relative MAPE improvement (evaluated as in (23)) across all test series for one-step-ahead forecasts. The distribution reveals that while the traditional model generally performs better at this horizon, the magnitude of improvement varies considerably across series. The median improvement (50th percentile) is approximately -7.5%, indicated by the middle vertical line, suggesting that for a typical series, the traditional model achieves modestly better one-step-ahead accuracy. The first quartile consists of 25% of the tested time series on which the AC-optimized model experienced equal to or greater than 25.4% degradation in terms of MAPE. The third quartile (75th percentile) is approximately -0.5%, indicating that the AC-optimized model showed practically no degradation or improvement for 25% of the test series.
Table 2. Model comparison using MAPE.
Statistic MLE-fitted MAPE AC-optimized MAPE % Improvement
25% Percentile 0.0025 0.0028 -25.4%
50% Percentile (Median) 0.0185 0.0321 -7.5%
75% Percentile 0.2055 0.2129 -0.5%
Figure 5. Percentage forecast MAPE improvement (%) of MLE-fitted model relative to AC-optimized model at one step ahead.
4.5. Sensitivity to Time Discounting Weights
We complement our discussion of weight functions in section 3.5 and assess their influence on optimizing multi-horizon accuracy and stability. We repeat the experiments using three additional weight types: uniform ($w(h)=1$), exponential ($w(h)=e^{-\frac{5}{24}h}$), and hyperbolic ($w(h)=\frac{1}{1+h}$), each normalized to sum up to 1.
Figure 6 presents the log mean vertical variance of all time series, sorted in ascending order and mapped to percentile. The mean vertical variance is defined in the same way as 22 and provides a comparison between the stability of forecast generated by models trained using each weight type. The log transformation is applied to adjust the scale. While all schemes improve stability, uniform weights produced the most stable forecasts at almost every percentile, followed by linear weights, then exponential and hyperbolic weights. This result is consistent with the heuristic that uniform weight assigns equal weights to short and long horizons, coercing the forecasts to stay stable. The other weights decay for longer horizon, with the rate of decay from slowest to fastest: linear, hyperbolic, exponential. Suppose one sets the rate of decay to approach infinity, which results in the weight converging to $w(h)=\delta(h=1)$ after normalization. This limit case is exactly equivalent to training the model based on one-step-ahead performance. Hence as the weight decays faster from uniform to exponential, the vertical variance increases and approaches what is learned by the one-step-ahead MLE model.
Figure 6. Effect of Weight on Optimizing Stability